# Process-Supervised LLM Recommenders via Flow-guided Tuning

### v1

@@ 3.3.3. Reward Setting
Remark: Many policy optimization methods employ complex process reward models (PRMs) (Lightman et al., 2024; Wang et al., 2024c) to enhance reasoning in large language models (Setlur et al., 2025; Zhang et al., 2024d), requiring significant computational resources for learning and verification. In contrast, our approach adopts heuristically assigned rewards, which are simple yet effective. This design avoids additional parameter learning and ensures efficient process supervision.
@@ 3.4. Fine-tuning LLMs through Process Rewards
This combined loss encourages the model to retain the supervised learning performance of SFT while leveraging the strengths of GFlowNets. Specifically, it incorporates the ability of GFlowNets to model diverse and proportional outcomes through process supervision, enabling the policy to generate items that not only align with the training data but also address issues of bias and fairness.
@@ 4.1.1. Dataset
We conduct experiments on three real-world datasets from Amazon review data [^2] https://cseweb.ucsd.edu/~jmcauley/datasets.html##\##amazon__\__reviews, including CDs, Video Games, and Movies. These datasets contain user review data from May 1996 to October 2018. Following the preprocessing strategy in (Bao et al., 2023a, b), we truncate the dataset based on time information, remove unpopular users and items with fewer than five interactions, and limit the maximum item sequence length to 10. The datasets are chronologically split into training, validation, and test sets in an 8:1:1 ratio, detailed statistical information is provided in Table 2.
@@ 4.1.2. Evaluation Protocal
For Accuracy, following prior work (Bao et al., 2023a, 2024b; Jiang et al., 2024), we adopt two widely used metrics: Hit Ratio (HR@K) and Normalized Discounted Cumulative Gain (NDCG@K). Fairness is measured using DGU@K and MGU@K, which quantify the discrepancy between the group distribution of recommended titles in the top-k results and their distribution in the training set’s historical sequences (Jiang et al., 2024). Titles are grouped by popularity: first, the frequency of each title in the training set is computed and sorted in descending order, then partitioned into eight equal-sized groups. Titles absent from the training set’s historical sequences are assigned to the least popular group. To assess Diversity, we use two metrics: (1) Entropy of Words (H), which calculates the entropy of all English words in the recommended titles, and (2) Type-Token Ratio (TTR), defined as the ratio of unique words to the total number of words in the recommendations.
@@ 4.1.3. Baseline
  • BIGRec (Bao et al., 2023a) is one of the earliest and most classic supervised fine-tuning (SFT) methods for directly generating item titles.
  • Temp (Bao et al., 2024b) adjusts the temperature coefficient for inference based on the trained BIGRec model.
  • $\boldsymbol{D^{3}}$ (Bao et al., 2024b) enhances BIGRec by removing length normalization to address amplification bias and incorporating a text-free assistant model, SASRec, in the inference stage to mitigate the homogeneity issue and improve recommendation diversity.
@@ 4.3. Next-item Recommendation Results (RQ2)
  • For methods specifically aimed at enhancing fairness, Flower outperforms IFairLRS across all metrics and datasets. This indicates that, compared to IFairLRS’s approach of reweighting the entire title, Flower’s use of probabilistic process supervision at the token level through GFlowNets achieves better training outcomes.
@@ 4.4. Flower as a Reference Policy (RQ3)
We select three DPO-based recommendation methods for comparison: DMPO (Bai et al., 2024) is framework that bridges the gap between LLMs and recommendation tasks by sampling multiple negative items as rejected responses. S-DPO (Chen et al., 2024) is method that incorporates multiple negative samples in user preference data and generalizes pairwise DPO loss to a softmax ranking loss. RosePO (Liao et al., 2024a) is a general framework that combines negative sampling strategies with personalized uncertainty to improve fairness, unbiasedness, and robustness. In our experiments, we sample rejected items based on item popularity distributions for comparison with RosePO. For RL-based methods, we implement PPO (Schulman et al., 2017), using the normalized SASRec score as the reward.
@@ 4.5.3. Performance Varying $\lambda$
The flow-guided loss and SFT loss can differ by 3 to 4 orders of magnitude, prompting us to explore the impact of $\lambda$ across different magnitudes. We report the results on the CDs dataset.
As shown in Fig. 6, as $\lambda$ increases from left to right, the influence of the SFT loss diminishes while that of the flow-guided loss increases. Accuracy, fairness, and diversity generally exhibit a trend of first improving and then declining, with the best performance observed around $\lambda=0.005$. When the weight of the flow-guided loss becomes excessively large, the lack of SFT loss constraints leads to a collapse in performance. Conversely, when $\lambda\leq 0.001$ (x-axis value of 10 or lower), the influence of the flow-guided loss becomes negligible, and the performance gradually converges to that of the SFT-only method.
@@ 5. Related Work
In this section, we briefly introduce LLMs for the recommendation task and the applications of GFlowNets.
@@ 5.1. LLMs for Recommendation
LLMs have demonstrated remarkable capabilities in text generation, reasoning, and generalization, prompting researchers to investigate their potential for personalized recommendation tasks. Recent studies have explored several paradigms for integrating LLMs into recommendation systems: (1) employing them as decision-making agents (Wang et al., 2024b, d; Shi et al., 2024; Wu et al., 2024a), (2) leveraging their vast knowledge to provide supplementary information that enhances the recommendation process (Sheng et al., 2024; Liu et al., 2024b; Geng et al., 2024; Harte et al., 2023b), and (3) utilizing them as user simulators to model user preferences and behaviors (Zhang et al., 2024b; Cai et al., 2024; Zhang et al., 2024a; Friedman et al., 2023; Wang et al., 2024e). As research evolved, supervised fine-tuning (SFT) emerged as a key method for adapting LLMs to domain-specific recommendation data, significantly improving their task performance (Liao et al., 2024b; Bao et al., 2023a; Liu et al., 2023b; Chen et al., 2025; Bao et al., 2023b; Zhang et al., 2023b). To further refine LLMs for individual user preferences and reduce biases, advanced post-training techniques, such as Direct Preference Optimization (DPO), have been proposed (Chen et al., 2024; Bai et al., 2024; Liao et al., 2024a; Gao et al., 2024). These approaches aim to align LLM outputs with human preferences using carefully constructed preference data.
However, as highlighted earlier, models fine-tuned with SFT often suffer from popularity bias. Specifically, in tasks that involve generating item titles, the models tend to favor items that are popular in the training set. This issue arises primarily from overfitting caused by the use of cross-entropy (CE) loss during fine-tuning. While some item tokenization methods have been introduced to represent items as unique identifier sequences (Rajput et al., 2024; Qu et al., 2024; Wang et al., 2024a; Yang et al., 2024a), these methods still rely on the CE loss paradigm, which perpetuates the overfitting problem. In this work, we propose a novel fine-tuning paradigm for recommendation systems that addresses the limitations of SFT.
@@ 5.2. Applications of GFlowNets
A key advantage of GFlowNets lies in their ability to sample diverse solutions while maintaining proportionality to the reward. Furthermore, GFlowNets demonstrate strong generalization capabilities, allowing them to handle states not encountered during training (Bengio et al., 2024a, b; Zhang et al., 2022; Malkin et al., 2024). These properties make GFlowNets particularly well-suited for tasks that require exploring a wide solution space, such as molecular design (Jain et al., 2022; Koziarski et al., 2024; Zhu et al., 2024) and structured prediction (Malkin et al., 2023; Deleu et al., 2022). In recommendation systems, GFlowNets have also shown significant potential. For example, Liu et al. (2023a) use GFlowNets to introduce diversity while maintaining quality in listwise recommendations, and Liu et al. (2024a) apply GFlowNets to enhance user retention while fostering exploration.
@@ 6. Conclusion
This work addresses key limitations of supervised fine-tuning (SFT) in LLM-based recommendation systems, namely limited diversity and amplified popularity bias, which undermine accuracy, fairness, and personalization. These issues stem from SFT’s reliance on cross-entropy loss, leading to overfitting and reinforcement of biases in the training data. To tackle these challenges, we propose Flower, a novel fine-tuning paradigm that leverages generative flow networks (GFlowNets) for process-level supervision. Flower models the recommendation task as a flow network, where item-level rewards, derived from item frequencies in the training data, are propagated to token-level rewards for next-token prediction. This approach aligns token probabilities with reward distributions, balancing accuracy with fairness and diversity in recommendations. Extensive experiments on three real-world sequential recommendation datasets demonstrate Flower’s effectiveness, achieving superior results in accuracy, diversity, and fairness compared to SFT. Moreover, applying alignment methods post-Flower tuning yields better performance than applying them to conventionally fine-tuned models.

### v2

@@ 3.3.3. Reward Setting
Remark: Many policy optimization methods employ complex process reward models (PRMs) (Lightman et al., 2024; Wang et al., 2024b) to enhance reasoning in large language models (Setlur et al., 2025a; Zhang et al., 2024), requiring significant computational resources for learning and verification. In contrast, our approach adopts heuristically assigned rewards, which are simple yet effective. This design avoids additional parameter learning and ensures efficient process supervision.
@@ 3.4. Fine-tuning LLMs through Process Rewards
This combined loss preserves the supervised performance of SFT while leveraging GFlowNets to promote diversity and reward-proportionality, enabling the policy to generate fairer and more representative items.
@@ 4.1.1. Dataset
We conduct experiments on three real-world datasets from Amazon review data [^2] https://cseweb.ucsd.edu/~jmcauley/datasets.html##\##amazon__\__reviews, including CDs, Video Games, and Movies. These datasets contain user review data from May 1996 to October 2018. Following the preprocessing strategy in (Bao et al., 2025), we truncate the dataset based on time information, remove unpopular users and items with fewer than five interactions, and limit the maximum item sequence length to 10. The datasets are chronologically split into training, validation, and test sets in an 8:1:1 ratio, detailed statistical information is provided in Table 2.
@@ 4.1.2. Evaluation Protocal
For Accuracy, following prior work (Bao et al., 2025, 2024; Jiang et al., 2024), we adopt two widely used metrics: Hit Ratio (HR@K) and Normalized Discounted Cumulative Gain (NDCG@K). Fairness is measured using DGU@K and MGU@K, which quantify the discrepancy between the group distribution of recommended titles in the top-k results and their distribution in the training set’s historical sequences (Jiang et al., 2024). Titles are grouped by popularity: first, the frequency of each title in the training set is computed and sorted in descending order, then partitioned into eight equal-sized groups. Titles absent from the training set’s historical sequences are assigned to the least popular group. To assess Diversity, we use two metrics: (1) Entropy of Words (H), which calculates the entropy of all English words in the recommended titles, and (2) Type-Token Ratio (TTR), defined as the ratio of unique words to the total number of words in the recommendations.
@@ 4.1.3. Baseline
  • BIGRec (Bao et al., 2025) is one of the earliest and most classic supervised fine-tuning (SFT) methods for directly generating item titles.
  • Temp (Bao et al., 2024) adjusts the temperature coefficient for inference based on the trained BIGRec model.
  • $\boldsymbol{D^{3}}$ (Bao et al., 2024) enhances BIGRec by removing length normalization to address amplification bias and incorporating a text-free assistant model, SASRec, in the inference stage to mitigate the homogeneity issue and improve recommendation diversity.
@@ 4.3. Next-item Recommendation Results (RQ2)
  • For methods specifically aimed at enhancing fairness, Flower outperforms IFairLRS across all metrics and datasets. This indicates that, compared to IFairLRS’s approach of reweighting the entire title, Flower’s use of probabilistic process supervision at the token level through GFlowNets achieves better results.
@@ 4.4. Flower as a Reference Policy (RQ3)
We select three DPO-based recommendation methods for comparison: DMPO (Bai et al., 2024) is framework that bridges the gap between LLMs and recommendation tasks by sampling multiple negative items as rejected responses. S-DPO (Chen et al., 2024) is method that incorporates multiple negative samples in user preference data and generalizes pairwise DPO loss to a softmax ranking loss. RosePO (Liao et al., 2024) is a general framework that combines negative sampling strategies with personalized uncertainty to improve fairness, unbiasedness, and robustness. In our experiments, we sample rejected items based on item popularity distributions for comparison with RosePO. For RL-based methods, we implement PPO (Schulman et al., 2017), using the frequency of each title in the training set as the reward.
@@ 4.5.3. Performance Varying $\lambda$
The flow-guided loss and SFT loss can differ by 3 to 4 orders of magnitude, prompting us to explore the impact of $\lambda$ across different magnitudes. We report the results on the CDs dataset. As shown in Fig. 6, as $\lambda$ increases from left to right, the influence of the SFT loss diminishes while that of the flow-guided loss increases. Accuracy, fairness, and diversity generally exhibit a trend of first improving and then declining, with the best performance observed around $\lambda=0.005$. When the weight of the flow-guided loss becomes excessively large, the lack of SFT loss constraints leads to a collapse in performance. Conversely, when $\lambda\leq 0.001$ (x-axis value of 10 or lower), the influence of the flow-guided loss becomes negligible, and the performance gradually converges to that of the SFT-only method.
@@ 5.1. LLMs for Recommendation
Large Language Models (LLMs) have shown strong capabilities in text generation, reasoning, and generalization, motivating their use in personalized recommendation tasks. Supervised fine-tuning (SFT) has become a core approach for adapting LLMs to domain-specific recommendation data, significantly boosting performance (Bao et al., 2025; Zhang et al., 2025a; Chen et al., 2025). To further align model outputs with user preferences and reduce bias, post-SFT training methods such as Direct Preference Optimization (DPO) have been proposed (Chen et al., 2024; Gao et al., 2025).
Despite these advances, SFT-based models often suffer from popularity bias, leading to filter bubbles and degraded user experience (Gao et al., 2023a, b). This is largely due to overfitting introduced by the cross-entropy loss used in fine-tuning, which biases the model toward frequently occurring items in the training set. While recent efforts represent items using unique identifier sequences (Rajput et al., 2024; Wang et al., 2024a), they remain within the CE loss framework and inherit its limitations.
In this work, we propose a novel fine-tuning paradigm that addresses the shortcomings of SFT and promotes more balanced, personalized recommendations.
5.2. Process Supervised
To better align LLMs with human preferences or enhance reasoning ability, a post-SFT alignment step is often introduced. Techniques like RLHF and DPO (Rafailov et al., 2023) apply outcome-level supervision by evaluating model responses holistically. However, such supervision is coarse-grained, making it hard to interpret or guide intermediate reasoning steps (Setlur et al., 2025b).
Recently, process supervision has gained attention for its ability to provide step-level feedback using Process Reward Models (PRMs) (Zhang et al., 2025b; Li and Li, 2025). Compared to outcome-level methods, process supervision offers more interpretable and direct optimization signals, leading to improved reasoning quality and alignment. Most existing approaches require learning parametric PRMs—e.g., modeling step-wise correctness probabilities (Wang et al., 2024b; Shao et al., 2024)—which is often impractical in recommendation tasks due to sparse user feedback. In this work, we adopt simple yet effective heuristic rewards to approximate step-level quality. This design avoids additional parameter learning and enables efficient process supervision in data-scarce settings.
@@ 5.3. Applications of GFlowNets
A key advantage of GFlowNets lies in their ability to sample diverse solutions while maintaining proportionality to the reward. Furthermore, GFlowNets demonstrate strong generalization capabilities, allowing them to handle states not encountered during training (Bengio et al., 2024a, b; Zhang et al., 2022; Malkin et al., 2022). These properties make GFlowNets particularly well-suited for tasks that require exploring a wide solution space, such as molecular design (Jain et al., 2022; Zhu et al., 2024) and structured prediction (Malkin et al., 2023). In recommendation systems, GFlowNets have also shown significant potential. For example, Liu et al. (2023) use GFlowNets to introduce diversity while maintaining quality in listwise recommendations, and Liu et al. (2024) apply GFlowNets to enhance user retention while fostering exploration.
@@ 6. Conclusion
This work addresses key limitations of supervised fine-tuning (SFT) in LLM-based recommendation systems, notably limited diversity and amplified popularity bias, which hinder accuracy, fairness, and personalization. These issues largely arise from the overfitting nature of cross-entropy loss used in SFT. To overcome these challenges, we propose Flower, a novel fine-tuning paradigm based on generative flow networks (GFlowNets) and process-level supervision. Flower frames recommendation as a flow network, propagating item-level rewards—derived from item frequencies—to token-level supervision. This aligns token generation with reward distributions, promoting balanced, diverse, and fair recommendations. Experiments on three real-world sequential recommendation datasets show that Flower outperforms SFT in accuracy, diversity, and fairness. Furthermore, applying alignment methods after Flower fine-tuning yields better results than applying them on top of standard SFT models.
@@
Acknowledgements.
This work is supported by the National Natural Science Foundation of China (62402470, U21B2026, 62121002, U24B20180), Anhui Provincial Natural Science Foundation (2408085QF189), the Fundamental Research Funds for the Central Universities of China (WK2100000053), and the advanced computing resources provided by the Supercomputing Center of the USTC.